OpenStudy (anonymous):

Evaluate: 10 − 64 ÷ 8 + 32 (earn a metal) fans too

OpenStudy (mertsj):

Do you know the order of operations?

OpenStudy (anonymous):

I'll do it for fun, but I think someone else deserves a metal.. 10-64=-54 8+32=40 -54=-6.75 so -6.75+40

OpenStudy (anonymous):

@Mertsj he's just asking questions

OpenStudy (mertsj):

ok. Thanks.

OpenStudy (anonymous):

You're welcome